OpenStudy (anonymous):

A culture started with 1,000 bacteria. After 5 hours, it grew to 1,200 bacteria. Predict how many bacteria will be present after 11 hours. Round your answer to a whole number

OpenStudy (anonymous):

N=b*g^t b=1000, 1200=1000*g^5 g=1.03713729 So: N=1000*1.03713729^t After 11 hours: N=1000*1.03713729^11 = 1 493.47771 So I'd say 1500.

OpenStudy (amistre64):

1000 * (e^((1 /5) * ln(12/10) * (11))) = 1,493.5